OpenStudy (anonymous):

Given that F(x) = x^ 2 + 2, evaluate F(1) + F(5).

OpenStudy (anonymous):

simple substitute x with 1 and 5

OpenStudy (anonymous):

F(1) = 1^ 2 + 2 F(5) = 5^ 2 + 2 F(1) + f(5)=?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

1^2 =?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1+2=?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

f(1) = 3 now f(5); what us 5^2=?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

25+2=?

OpenStudy (anonymous):

o i got it

OpenStudy (anonymous):

let me know the final answer and i will tell u if it is correct